OpenStudy (anonymous):

Simplify. 3 square root 5 - 2 square root 7 + square root 45 - square root 28 (1 point)

OpenStudy (anonymous):

\[3\sqrt{5} - 2\sqrt{7} + \sqrt{45} - \sqrt{28}\]